OpenStudy (anonymous):

1.Solve the inequality. Show your work. |r + 3| ≥ 7

OpenStudy (anonymous):

|r + 3| >= 7 r + 3 >= 7 or r + 3 <= -7 r >= 4 or r <= -10

OpenStudy (anonymous):

|r + 3| ≥ 7 r+3≥7 r≥7-3 r=4= -10

OpenStudy (anonymous):

just transpose only the number in the left so u can get the value of r.